Circular No. 6379

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1996W IN NGC 4027 W. D. Li, Y. L. Qiu, Q. Y. Qiao, J. Ma, and J. Y. Hu, Beijing As-

tronomical Observatory (BAO), report the discovery of a supernova found on CCD images taken on Apr. 10 and 11 with the 0.60-m reflector during the BAO supernova survey. SN 1996W (V ∼ 16) is located 3′′ west and 40′′ north of the center of NGC 4027 (α = 11^{h}59^{m}30^{s}.1, δ = −19^{o}16′05′′, equinox 2000.0). The object is not present on an image taken on Mar. 1.

L. Wang, and J. C. Wheeler, University of Texas at Austin and Mc-Donald Observatory, report: “SN 1996W was observed on Apr. 12 UT at the McDonald Observatory 2.1-m telescope (+ Imaging Grism Instrument). Inspections of a fully-reduced spectrum (range 424.7–813.5 nm, resolution 2.0 nm) show that this is a type-II supernova prior to optical maximum. Weak but conspicuous Hα and Hβ P-Cyg profiles are superimposed on a very blue continuum. The redshift of the host galaxy is ∼ 1300 km/s as derived from narrow emission lines of the neighboring H II region. The expansion velocity of the supernova ejecta is ∼ 14 300 km/s, as derived from the Hα P-Cyg profile. The He I/Na I 587.6-nm feature is also present.”

SUPERNOVA 1996V IN NGC 3644 P. Sicoli reports the following precise position of SN 1996V obtained

at Sormano by V. Giuliani and P. Ghezzi with the 0.5-m reflector (+ CCD) on Apr. 11.939 UT: α = 11^{h}21^{m}31^{s}.23, δ = +2^{o}48′40′′.4 (equinox 2000.0).

COMET 48P/JOHNSON CCD R magnitude estimates by W. Offutt, Cloudcroft, NM (0.6-m

reflector): Mar. 16.32 UT, 20.9; Apr. 9.23, 20.7.

COMET 116P/WILD 4 Total visual magnitude estimates: Mar. 8.83 UT, 12.5 (K. Hornoch,

Lelekovice, Czech Republic, 0.35-m reflector); 10.92, 12.2 (T. Vanmunster, Landen, Belgium, 0.35-m reflector); 19.81, 12.1 (V. Znojil, Brno, Czech Republic, 25×100 binoculars); 24.04, 12.4 (J. M. San Juan, Teruel, Spain, 0.44-m reflector); Apr. 8.92, 12.6 (J. Carvajal, Madrid, Spain, 0.21-m reflector); 11.06, 12.1 (G. W. Kronk, Troy, IL, 33.3-cm reflector).

1996 April 12  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green